# BOTS: A Unified Framework for Bayesian Online Task Selection in LLM Reinforcement Finetuning

### v1

@@ Appendix C More Related Works
Reinforcement finetuning (RFT) has become a pivotal paradigm for aligning Large Language Models (LLMs) with human preferences and enhancing reasoning capabilities. Early approaches, commonly known as Reinforcement Learning from Human Feedback (RLHF) (Ouyang et al., 2022), train a reward model to approximate human preferences and then optimize the policy using reinforcement learning algorithms such as PPO (Schulman et al., 2017). Subsequent research introduced direct preference optimization methods such as DPO (Rafailov et al., 2023) and its variants KTO (Ethayarajh et al., 2024), ORPO (Hong et al., 2024), and SimPO (Meng et al., 2024), which bypass the need to train an explicit reward model. More recently, researchers have observed that when rewards are automatically verifiable—such as in mathematics or code generation—RFT can substantially enhance reasoning performance, leading to the emergence of verifiable-RL frameworks like RLVR, where GRPO (Shao et al., 2024) serves as a PPO-style adaptation designed for such settings. Building on these foundations, variants like DAPO (Yu et al., 2025) and GSPO (Zheng et al., 2025) further refine the optimization dynamics, while other studies explore entropy regularization (Cui et al., 2025; Wang et al., 2025a) and reward shaping (Chen et al., 2025a; Pan et al., 2025). Despite these algorithmic advances, a critical bottleneck remains—the data curriculum, i.e., how tasks are selected and presented during RFT, which crucially determines training efficiency.
@@
The importance of data difficulty in RFT is increasingly recognized, leading to the creation of highly challenging benchmarks (Albalak et al., 2025; He et al., 2025b; Gao et al., 2025). Beyond dataset creation, data selection and curriculum design has also become a central topic of investigation. For Supervised Finetuning (SFT), a variety of data selection and curriculum strategies have been widely studied based on quality (Liu et al., 2024; Li et al., 2024a), diversity (Ling et al., 2025; Lu et al., 2024), or pre-assessed difficulty (Xu et al., 2025; Ye et al., 2025). However, such static methods are ill-suited for the dynamic nature of RFT. This has motivated a focus on dynamic curricula that adapt to the model’s evolving capabilities, typically by leveraging the notion of task difficulty (Cheng et al., 2025; Pikus et al., 2025; Li et al., 2025b).
@@
Existing task selection strategies for RFT vary in their approach to leveraging task difficulty. An early line of work, directly inspired by curriculum learning, employs offline curricula that schedule tasks along a fixed easy-to-hard trajectory (Parashar et al., 2025; Shen et al., 2025; Zhu et al., 2025; Wen et al., 2025; Li et al., 2025a). While simple and intuitive, these methods are non-adaptive and cannot respond to the model’s real-time learning progress. To address this, online selection strategies have emerged. A straightforward approach is sampling-based task filtering, which uses extra rollouts to evaluate and discard tasks that are too easy or too hard, thereby incurring significant computational overhead (Yu et al., 2025; Bae et al., 2025). To avoid this cost, recent works attempt to predict task success rates. Some frame task selection as a non-stationary multi-armed bandit (MAB) problem, but often treat tasks as independent arms, thus overlooking cross-task relationships (Chen et al., 2025b; Qu et al., 2025). Others use a small reference set to predict the difficulty of other tasks via similarity kernels, but this still requires extra rollouts and discards valuable historical information (Sun et al., 2025; Wang et al., 2025b).
@@ E.2 Attention-Based Adaptive Difficulty Estimate
For example, Sun et al. (2025) proposed to use an attention-like kernel structure:
@@ E.3 Interpolation-Based Implicit Evidence
Unlike kernel-based estimators (Sun et al., 2025) (see Appendix E.2 for a brief introduction), which are confined to the convex hull of observed tasks, our interpolation-based estimator naturally supports extrapolation. For example, if the reference pair consists of a strong and a stronger model, the capability coefficient $\mu_{t}(\mathcal{B}_{t})$ defined in Section 3.3 may fall outside $[0,1]$, yielding extrapolated predictions. The clipping step in Equation (5) ensures these estimates remain within the feasible range $[0,1]$.
@@
The effectiveness of the implicit evidence mechanism depends critically on the choice of weak and strong reference models, which should ideally bracket the expected capability trajectory of the model under training. If the training model is substantially stronger than the strong reference or weaker than the weak reference, our method must rely on extrapolation, whose accuracy may degrade. In practice, we find that standard publicly available checkpoints—Qwen2.5-7B-Instruct (Yang et al., 2024) and Qwen3-32B (Yang et al., 2025)—provide a robust and practical basis, even when training significantly smaller models such as Qwen2.5-1.5B-Instruct (Yang et al., 2024). A systematic ablation on the choice of reference models remains an important direction for future work.
@@ E.4.3 Alternative Plug-In for Implicit Evidence
Our interpolation-based estimator provides an extremely lightweight way to generate implicit evidence without additional rollouts, but it is not the only possible choice. More expressive alternatives could be explored, such as kernel-based predictors (Sun et al., 2025), task-embedding regressors, or small auxiliary models trained jointly with the main model. These alternatives may improve predictive accuracy, especially when the reference models poorly bracket the training model’s capability. However, they also introduce a trade-off: richer implicit evidence often requires higher computational and storage costs. Systematically characterizing this trade-off—between accuracy and efficiency in implicit evidence—remains an open research question and a promising direction for future work.

### v2

@@ Appendix C More Related Works
Reinforcement finetuning (RFT) has become a pivotal paradigm for aligning Large Language Models (LLMs) with human preferences and enhancing reasoning capabilities. Early approaches, commonly known as Reinforcement Learning from Human Feedback (RLHF) (ouyang2022rlhf), train a reward model to approximate human preferences and then optimize the policy using reinforcement learning algorithms such as PPO (2017ppo). Subsequent research introduced direct preference optimization methods such as DPO (2023dpo) and its variants KTO (2024kto), ORPO (2024orpo), and SimPO (2024simpo), which bypass the need to train an explicit reward model. More recently, researchers have observed that when rewards are automatically verifiable—such as in mathematics or code generation—RFT can substantially enhance reasoning performance, leading to the emergence of verifiable-RL frameworks like RLVR, where GRPO (grpo2024) serves as a PPO-style adaptation designed for such settings. Building on these foundations, variants like DAPO (yu2025dapo) and GSPO (2025gspo) further refine the optimization dynamics, while other studies explore entropy regularization (2025primeentropy; 2025entropy28) and reward shaping (2025seedgrpo; 2025trinity). Despite these algorithmic advances, a critical bottleneck remains—the data curriculum, i.e., how tasks are selected and presented during RFT, which crucially determines training efficiency.
@@
The importance of data difficulty in RFT is increasingly recognized, leading to the creation of highly challenging benchmarks (2025bigmath; he2025deepmath; gao2025omni). Beyond dataset creation, data selection and curriculum design has also become a central topic of investigation. For Supervised Finetuning (SFT), a variety of data selection and curriculum strategies have been widely studied based on quality (2024deita; 2024superfilter), diversity (2025daar; 2024instag), or pre-assessed difficulty (2025mindgym; 2025limo). However, such static methods are ill-suited for the dynamic nature of RFT. This has motivated a focus on dynamic curricula that adapt to the model’s evolving capabilities, typically by leveraging the notion of task difficulty (cheng2025revisiting; 2025hardexamples; li2025can).
@@
Existing task selection strategies for RFT vary in their approach to leveraging task difficulty. An early line of work, directly inspired by curriculum learning, employs offline curricula that schedule tasks along a fixed easy-to-hard trajectory (parashar2025curriculum; shen2025skywork; zhu2025eduflow; wen2025structured; li2025temple). While simple and intuitive, these methods are non-adaptive and cannot respond to the model’s real-time learning progress. To address this, online selection strategies have emerged. A straightforward approach is sampling-based task filtering, which uses extra rollouts to evaluate and discard tasks that are too easy or too hard, thereby incurring significant computational overhead (yu2025dapo; bae2025online). To avoid this cost, recent works attempt to predict task success rates. Some frame task selection as a non-stationary multi-armed bandit (MAB) problem, but often treat tasks as independent arms, thus overlooking cross-task relationships (chen2025self; qu2025can). Others use a small reference set to predict the difficulty of other tasks via similarity kernels, but this still requires extra rollouts and discards valuable historical information (sun2025improving; wang2025dump).
@@ E.2 Attention-Based Adaptive Difficulty Estimate
For example, sun2025improving proposed to use an attention-like kernel structure:
@@ E.3 Interpolation-Based Implicit Evidence
Unlike kernel-based estimators (sun2025improving) (see Appendix E.2 for a brief introduction), which are confined to the convex hull of observed tasks, our interpolation-based estimator naturally supports extrapolation. For example, if the reference pair consists of a strong and a stronger model, the capability coefficient $\mu_{t}(\mathcal{B}_{t})$ defined in Section 3.3 may fall outside $[0,1]$, yielding extrapolated predictions. The clipping step in Equation (5) ensures these estimates remain within the feasible range $[0,1]$.
@@
The effectiveness of the implicit evidence mechanism depends critically on the choice of weak and strong reference models, which should ideally bracket the expected capability trajectory of the model under training. If the training model is substantially stronger than the strong reference or weaker than the weak reference, our method must rely on extrapolation, whose accuracy may degrade. In practice, we find that standard publicly available checkpoints—Qwen2.5-7B-Instruct (Yang2024Qwen25TR) and Qwen3-32B (yang2025qwen3)—provide a robust and practical basis, even when training significantly smaller models such as Qwen2.5-1.5B-Instruct (Yang2024Qwen25TR). A systematic ablation on the choice of reference models remains an important direction for future work.
@@ E.4.3 Alternative Plug-In for Implicit Evidence
Our interpolation-based estimator provides an extremely lightweight way to generate implicit evidence without additional rollouts, but it is not the only possible choice. More expressive alternatives could be explored, such as kernel-based predictors (sun2025improving), task-embedding regressors, or small auxiliary models trained jointly with the main model. These alternatives may improve predictive accuracy, especially when the reference models poorly bracket the training model’s capability. However, they also introduce a trade-off: richer implicit evidence often requires higher computational and storage costs. Systematically characterizing this trade-off—between accuracy and efficiency in implicit evidence—remains an open research question and a promising direction for future work.